The speed of sound in a gas, in which two waves of wavelength 1.0 m and 1.02 m produce 6 beats per second, is approximately :
 

Grade:12th pass

2 Answers

Arun
4 years ago
Dear Neha
 
As we know
 
1.02V - 1.00V=6 x 1.02
0.02 V = 6.12
V = 306 m/sec
 
Hope it helps
 
 
Thanks. & Regards
Arun
Vikas TU
4 years ago
Dear student 
f1 = v/lambda 1= V/1.00
f2 = V/Lambda2 
f1 -f2 =  6 
0.02V = 6.06
V = 6.06 / 0.02 = 303 m/s
Good Luck